The distance between the center point of 2 consecutive posts in a fence is L ft. The posts are each x in. wide. The n pickets are each a in. wide equally spaced. Determine width of each space.
Not sure how to continue.

Answer by rothauserc(4718)

Draw a picture of the problem
:
L = 2(x/2) + na + (n+1)s where s the width of each space
:
L = x + na + (n+1)s
:
(n+1)s = L - x - na
:
**********************
s = (L -x -na) / (n+1)
**********************
: